OpenStudy (anonymous):

If p(q-r)x^2 +q(r-p)x + r(p-q) has equal roots,then, 2/q=?? a)p+ 1/r c)p+r b) 1/p +r d) 1/p + 1/r

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

\[D = b^2 - 4ac\] For equal roots D = 0

OpenStudy (anonymous):

i know that, tell me what is 2/q and how ??

OpenStudy (anonymous):

cosidering symmtry of that qn ans wud be d :P

OpenStudy (anonymous):

It is the value you have to find from this approach...

OpenStudy (anonymous):

just put those things in nd solve....u l get it

OpenStudy (anonymous):

got your answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i have solved it till pqr(p-q)-pr^2(p-q)=0

OpenStudy (anonymous):

wait ill check

OpenStudy (anonymous):

i am sorry i am not getting answer now will tell you in the evening

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

any other problem in quadratic?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you show me how have you arrived at that equation??

OpenStudy (anonymous):

just simplified everything.

OpenStudy (anonymous):

I want to see if you want to show me..

OpenStudy (anonymous):

\[q^{2}r^{2}-2pq^{2}r+p^{2}r^{2}-4pq(pr-qr)-pr(pr-qr)=0\]

OpenStudy (anonymous):

this came by putting D=0.

OpenStudy (anonymous):

How you have put the value of 4ac??

OpenStudy (anonymous):

\[b^2 - 4ac = 0\] \[q^2(r-p)^2 - 4pr(p-q)(q-r) = 0\]

OpenStudy (anonymous):

yes.

OpenStudy (phi):

\[ q^2(r-p)^2-4pr(q-r)(p-q)=0\] make this a quadratic of \( q^2 \) First, expand the 2nd term to get \[ -4pr(pq-pr-q^2+qr)= -4p^2rq+4p^2r^2+4pq^2r-4pqr^2 \] \[= 4pr q^2 -4pr(p+r) q +4p^2r^2 \] this will make the expression \[ ((r-p)^2+4pr) q^2 -4pr(p+r) q + 4p^2r^2=0 \] the coefficient of q^2 can be simplified to (p+r)^2 \[ (p+r)^2 q^2 -4pr(p+r) q + 4p^2r^2=0 \] solve for q using the quadratic formula \[ \frac{4pr(p+r)}{2(p+r)^2} \pm \frac{\sqrt{D}}{2(p+r)^2} \] where D is \[ 16p^2r^2(p+r)^2 - 16(p+r)^2 p^2r^2=0 \] we find q \[ q= \frac{2pr}{(p+r)} \] and \[ \frac{2}{q}= \frac{p+r}{pr}= \frac{1}{r}+\frac{1}{p} \]